### v1

# P3P Made Easy

Abstract
We present a novel algebraic solution to the Perspective-Three-Point (P3P) problem, which aims to recover the absolute pose of a calibrated camera from three 2D–3D correspondences. Our method reformulates the problem into a quartic polynomial with coefficients that are analytically simple and computationally efficient. Despite its simplicity, the proposed solver achieves accuracy and runtime performance comparable to state-of-the-art methods. Extensive experiments on synthetic datasets validate its robustness and efficiency. This combination of simplicity and performance makes our solver appealing for both real-time systems and educational contexts, where interpretability and reliability are critical.
1 Introduction
Estimating the pose (i.e., position and orientation) of a calibrated camera from image measurements is a foundational problem in computer vision. It is relevant to a wide variety of applications, including 3D reconstruction [1], Structure-from-Motion (SfM) [19], augmented reality (AR) [14], visual odometry [20], visual localization [18] and simultaneous localization and mapping (SLAM) [13]. A particularly important and well-studied variant is the Perspective-Three-Point (P3P) problem, where the absolute pose of a calibrated camera is determined from exactly three 2D-to-3D point correspondences (see Fig. 1).
The importance of P3P arises not only from its theoretical elegance but also from its practical utility. In real-world systems, rapid and reliable pose estimation is critical. In AR, for instance, aligning virtual content with the physical world requires accurate and low-latency camera pose estimation. In SLAM and visual(-inertial) navigation, the camera pose must be tracked accurately from sparse and noisy data. In such contexts, the ability to estimate pose from only three correspondences makes P3P a valuable tool, particularly for minimal-solver-based pipelines where efficiency and robustness are crucial.
[…]
Over the years, various P3P methods have been proposed, and they are typically classified by the degree and form of the polynomial equations they produce — most commonly quartic or cubic. While quartic-based solvers have been traditionally favored for their analytical simplicity, cubic-based solvers have recently attracted increasing interest due to their numerical stability and computational efficiency [17, 4].
In this work, we propose a novel quartic-based P3P solver that is numerically stable and computationally efficient. Our main contribution is twofold:
  1. 1.
    We show that the P3P problem can be reduced to solving a quartic polynomial with coefficients that are remarkably straightforward to derive and interpret. The derivation relies only on elementary mathematics, making the approach accessible even at the high school level. To the best of our knowledge, this degree of simplicity has not been previously demonstrated in the literature on the P3P problem.
  2. 2.
    We propose a set of practical implementation strategies, which are shown to significantly enhance the numerical accuracy of the solution.
While many existing P3P solvers derive quartic or cubic polynomials using complex symbolic or geometric tools, our method stands out for its exceptional simplicity. We arrive at the quartic using only elementary algebraic reasoning, without relying on coordinate reparameterization [12, 15] or intermediate coordinate systems [2, 16]. This leads to a solver that is not only efficient and numerically stable but also easy to derive, implement, and verify, making it particularly well-suited for both practical deployment and instructional purposes.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 formally defines the P3P problem, and Section 4 introduces the proposed method. Evaluation results and analysis are presented in Sections 5 and 6, respectively. Finally, Section 7 concludes the paper.
2 Related Work
The P3P problem has a long and rich history. The first explicit treatment of this problem appeared in a paper by Grunert in 1841 [9]. Grunert formulated the task of determining camera pose from three spatial correspondences and showed that it could yield up to four real solutions. His approach relied on geometric constraints and trigonometric relationships, and it became the foundation for early photogrammetric methods (see the survey by Haralick et al. [10]). Building on Grunert’s ideas, in 1937, Finsterwalder and Scheufele [5] proposed a formulation that effectively reduced the P3P problem to a cubic equation, making their method the first cubic-based P3P solver. The practical impact of P3P solvers was amplified by the introduction of RANSAC in 1981 by Fischler and Bolles [6], where they also introduced one of the earliest closed-form quartic-based solvers for P3P. In 1991, Haralick et al. [10] offered one of the earliest comprehensive overviews of P3P methods, bringing together and organizing various geometric and algebraic approaches developed over previous decades.
Based on the degree of the polynomial equation they solve, existing P3P solvers can be categorized into two main groups: quartic-based and cubic-based. Below, we summarize key methods in each group.
Quartic-based P3P solvers: Quartic-based solvers have been the most common historically. Gao et al. [7] applied Wu-Ritt’s zero decomposition technique [22] to eliminate variables and derive a quartic polynomial whose real roots capture the geometric constraints of the P3P problem. Kneip et al. [12] proposed a parameterization that directly computes the camera pose without estimating 3D points in the camera frame, which reduced the problem to solving a quartic polynomial using two parameters, avoiding the singular value decomposition. Masselli and Zell [15] proposed an alternative prameterization using different geometric constraints. Ke and Roumeliotis [11] systematically eliminated the camera position and feature distances, solving directly for orientation by deriving a quartic polynomial in a single rotation parameter. Banno [2] introduced an algebraic approach that represents the camera pose parameters as a linear combination of null-space vectors derived from the P3P constraints. Nakano [16] improved on Banno’s method by introducing a more efficient coordinate setup, which reduced algebraic complexity and computational cost. Recently, Wu et al. [23] proposed a conic transformation approach to the P3P problem that maps one of the conics into a canonical parabola, simplifying the intersection computation and eliminating the need for complex arithmetic.
Cubic-based P3P solvers: A smaller but notable group of methods reduces the P3P problem to solving a cubic polynomial. These methods generally aim to reduce the number of candidate solutions and improve numerical stability by simplifying the algebraic structure. Early examples of this type include the works of Finsterwalder and Scheufele [5] and Grafarend et al. [8]. A more recent example is the cubic-based solver proposed by Persson and Nordberg [17], where the underlying elliptic equations are solved by diagonalization that requires a single real root of a cubic polynomial. In [4], Ding et al. revisit the cubic-based approach and systematically characterize all eight possible conic intersection cases. By explicitly handling each case, their method improves both numerical stability and computational efficiency over prior work.
3 Problem Definition
Consider a calibrated camera located at the origin, observing three 3D points. Let $\mathbf{m}_{1}$, $\mathbf{m}_{2}$ and $\mathbf{m}_{3}$ be the unit bearing vectors directed toward the corresponding 3D points $\mathbf{X}_{1}$, $\mathbf{X}_{2}$ and $\mathbf{X}_{3}$, as shown in Fig. 1. Let $\mathbf{R}$ and $\mathbf{t}$ be the unknown rotation matrix and translation vector applied to the 3D points, such that
$d_{1}\mathbf{m}_{1}$ $=\mathbf{RX}_{1}+\mathbf{t},$ (1)
$d_{2}\mathbf{m}_{2}$ $=\mathbf{RX}_{2}+\mathbf{t},$ (2)
$d_{3}\mathbf{m}_{3}$ $=\mathbf{RX}_{3}+\mathbf{t},$ (3)
where $d_{1}$, $d_{2}$ and $d_{3}$ are the unknown depths of the points. Our goal is to find $\mathbf{R}$ and $\mathbf{t}$ (and optionally $\mathbf{d}_{1}$, $\mathbf{d}_{2}$ and $\mathbf{d}_{3}$) that satisfy (1)–(3), given the unit vectors $\left(\mathbf{m}_{1},\mathbf{m}_{2},\mathbf{m_{3}}\right)$ and the corresponding 3D points $\left(\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\right)$ as input.
Figure 1: An illustration of the P3P problem. (a) Ground truth: The unit bearing vectors from the camera center pass through the corresponding 3D points in space. (b) Problem input: An unknown rigid-body transformation has been applied to the given 3D points. As a result, the bearing vectors do not pass through the 3D points. Our goal is to find the unknown rigid-body transformation, given these bearing vectors and 3D points as input.
4 Proposed Method
4.1 Theory
We define the following variables:
$m_{12}$ $=\mathbf{m}_{1}^{\top}\mathbf{m}_{2},$ (4)
$m_{13}$ $=\mathbf{m}_{1}^{\top}\mathbf{m}_{3},$ (5)
$m_{23}$ $=\mathbf{m}_{2}^{\top}\mathbf{m}_{3},$ (6)
$s_{12}$ $=\lVert\mathbf{X}_{1}-\mathbf{X}_{2}\rVert^{2},$ (7)
$s_{13}$ $=\lVert\mathbf{X}_{1}-\mathbf{X}_{3}\rVert^{2},$ (8)
$s_{23}$ $=\lVert\mathbf{X}_{2}-\mathbf{X}_{3}\rVert^{2}.$ (9)
Note that since $\mathbf{m}_{1}$, $\mathbf{m}_{2}$ and $\mathbf{m}_{3}$ are unit vectors, $m_{12}$, $m_{13}$ and $m_{23}$ represent the cosines of the angles between each pair of bearing vectors (see the Fig. 1a). According to the law of cosines, we have
$d_{1}^{2}+d_{2}^{2}-2d_{1}d_{2}m_{12}$ $=s_{12},$ (10)
$d_{1}^{2}+d_{3}^{2}-2d_{1}d_{3}m_{13}$ $=s_{13},$ (11)
$d_{2}^{2}+d_{3}^{2}-2d_{2}d_{3}m_{23}$ $=s_{23}.$ (12)
Here, we introduce the following variables:
$x=\frac{d_{1}}{d_{3}},\quad y=\frac{d_{2}}{d_{3}}.$ (13)
Substituting $d_{1}=xd_{3}$ and $d_{2}=yd_{3}$ into (10)–(12) and rearranging them, we obtain the following equations:
$d_{3}^{2}$ $=\frac{s_{12}}{x^{2}-2xym_{12}+y^{2}},$ (14)
$d_{3}^{2}$ $=\frac{s_{13}}{x^{2}-2xm_{13}+1},$ (15)
$d_{3}^{2}$ $=\frac{s_{23}}{y^{2}-2ym_{23}+1}.$ (16)
Equating the right-hand side of (16) with that of (14) and (15), respectively, we get the following two equations:
$x^{2}-2m_{12}xy+(1-a)y^{2}+2am_{23}y-a=0,$ (17)
$x^{2}-by^{2}-2m_{13}x+2bm_{23}y+1-b=0,$ (18)
where
$a=\frac{s_{12}}{s_{23}},\quad b=\frac{s_{13}}{s_{23}}.$ (19)
Note that (17) and (18) are the same set of equations used in [4, 23] to derive a polynomial equation in $x$. From here on, however, we take a much simpler approach than these two works: First, we rearrange (18) as
$y^{2}=\frac{x^{2}-2m_{13}x+2bm_{23}y+1-b}{b}.$ (20)
Substituting this into (17), we obtain an equation involving only $x^{2}$, $x$ and $y$, which we then simplify to isolate $y$ as a rational function of $x$:
$y=\frac{(-a+b+1)x^{2}+2(a-1)m_{13}x-a-b+1}{2b(m_{12}x-m_{23})}.$ (21)
Using the definitions in (19), this can also be written as
$y=\frac{Ax^{2}+Bx+C}{2s_{13}(m_{12}x-m_{23})},$ (22)
where
$A$ $=-s_{12}+s_{23}+s_{13},$ (23)
$B$ $=2(s_{12}-s_{23})m_{13},$ (24)
$C$ $=-s_{12}+s_{23}-s_{13}.$ (25)
Plugging (22) into (18) yields quartic polynomial in $x$. Conveniently, it turns out that all of the coefficients of this polynomial contain a common factor, $s_{13}/s_{23}$, which is $b$ from (19). Factoring this out, we get:
$c_{4}x^{4}+c_{3}x^{3}+c_{2}x^{2}+c_{1}x+c_{0}=0,$ (26)
where
$\begin{split}c_{4}={}&-s_{12}^{2}+2s_{12}s_{13}+2s_{12}s_{23}-s_{13}^{2}\\ &+4s_{13}s_{23}m_{12}^{2}-2s_{13}s_{23}-s_{23}^{2},\end{split}$ (27)
$\begin{split}c_{3}={}&4s_{12}^{2}m_{13}-4s_{12}s_{13}m_{12}m_{23}-4s_{12}s_{13}m_{13}\\ &-8s_{12}s_{23}m_{13}+4s_{13}^{2}m_{12}m_{23}\\ &-8s_{13}s_{23}m_{12}^{2}m_{13}-4s_{13}s_{23}m_{12}m_{23}\\ &+4s_{13}s_{23}m_{13}+4s_{23}^{2}m_{13},\end{split}$ (28)
$\begin{split}c_{2}={}&-4s_{12}^{2}m_{13}^{2}-2s_{12}^{2}+8s_{12}s_{13}m_{12}m_{13}m_{23}\\ &+4s_{12}s_{13}m_{23}^{2}+8s_{12}s_{23}m_{13}^{2}+4s_{12}s_{23}\\ &-4s_{13}^{2}m_{12}^{2}-4s_{13}^{2}m_{23}^{2}+2s_{13}^{2}\\ &+4s_{13}s_{23}m_{12}^{2}+8s_{13}s_{23}m_{12}m_{13}m_{23}\\ &-4s_{23}^{2}m_{13}^{2}-2s_{23}^{2},\end{split}$ (29)
$\begin{split}c_{1}={}&4s_{12}^{2}m_{13}-4s_{12}s_{13}m_{12}m_{23}\\ &-8s_{12}s_{13}m_{13}m_{23}^{2}+4s_{12}s_{13}m_{13}\\ &-8s_{12}s_{23}m_{13}+4s_{13}^{2}m_{12}m_{23}\\ &-4s_{13}s_{23}m_{12}m_{23}-4s_{13}s_{23}m_{13}\\ &+4s_{23}^{2}m_{13},\end{split}$ (30)
$\begin{split}c_{0}={}&-s_{12}^{2}+4s_{12}s_{13}m_{23}^{2}-2s_{12}s_{13}\\ &+2s_{12}s_{23}-s_{13}^{2}+2s_{13}s_{23}-s_{23}^{2}.\end{split}$ (31)
Solving (26) yields up to four real solutions for $x$. Plugging these $x$ values into (22) gives the corresponding $y$ values. By considering only the positive real $x$ and $y$ values, we can find $d_{3}$ using either (14), (15) or (16). Subsequently, $d_{1}$ and $d_{2}$ can be found using (13). Once all depths are found, we refine them using the Gauss-Newton method on (10)–(12). Afterwards, we align the set of 3D points $\left(d_{1}\mathbf{m}_{1},d_{2}\mathbf{m}_{2},d_{3}\mathbf{m}_{3}\right)$ to $\left(\mathbf{X}_{1},\mathbf{X}_{2},\mathbf{X}_{3}\right)$ using the method proposed in [17]:
$\mathbf{R}$ $=\begin{bmatrix}\mathbf{Y}_{1},\mathbf{Y}_{2},\mathbf{Y}_{1}\times\mathbf{Y}_{2}\end{bmatrix}\mathbf{X}^{-1},$ (32)
$\mathbf{t}$ $=d_{1}\mathbf{m}_{1}-\mathbf{R}\mathbf{X}_{1},$ (33)
where
$\mathbf{Y}_{1}=d_{1}\mathbf{m}_{1}-d_{2}\mathbf{m}_{2},\quad\mathbf{Y}_{2}=d_{1}\mathbf{m}_{1}-d_{3}\mathbf{m}_{3},$ (34)
$\mathbf{X}=\begin{bmatrix}\mathbf{X}_{1}-\mathbf{X}_{2},\mathbf{X}_{1}-\mathbf{X}_{3},(\mathbf{X}_{1}-\mathbf{X}_{2})\times(\mathbf{X}_{1}-\mathbf{X}_{3})\end{bmatrix}.$ (35)
As a result, we obtain up to four valid solutions for $(\mathbf{R},\mathbf{t})$, corresponding to the real roots of the quartic.
4.2 Implementation details
We found that several subtle implementation choices can have a substantial impact on numerical precision. When following the instructions given in the previous paragraph, there are multiple ways, all mathematically equivalent, to go from (26) to the final solution $(\mathbf{R},\mathbf{t})$. Empirically, we found that some methods are numerically more precise than the others. In this section, we report our findings on the numerical impact of some seemingly minor choices. Revealing the exact mechanism underlying these heuristics is beyond the scope of this work and is left for future work.
  1. 1.
    Solving a quartic polynomial: To solve (26), we use two different (but mathematically equivalent) versions of Ferrari’s method: (1) Ferrari-Lagrange method [21] [^1] It solves a monic quartic polynomial $x^{4}+ax^{3}+bx^{2}+cx+d=0$ using the cubic solvent $y^{3}-by^{2}+(ac-4d)y+(4bd-a^{2}d-c^{2})=0$., and (2) Classical Ferrari method [3] [^2] It solves a depressed quartic polynomial $u^{4}+au^{2}+bu+c=0$ using the cubic solvent $8y^{3}+20ay^{2}+(16a^{2}-8c)y+(4a^{3}-4ac-b^{2})=0$.. A similar approach is proposed in [23], where Method 1 is used when the leading coefficient has magnitude less than $10^{4}$, and Method 2 is used when Method 1 could not find any real root or when the leading coefficient has magnitude larger than $10^{4}$. In this work, however, we use a different criterion: We use Method 1 when $\left|\frac{c_{3}}{c_{4}}\right|>10$; otherwise, we use Method 2. In both Method 1 and Method 2, we discard a root if its computation involves a negative argument under a square root, as in [23], to ensure that all intermediate computations remain real-valued. For the detailed description of the two methods, see the supplementary material of [23].
  2. 2.
    Reindexing the points: We found that indexing the points in a certain order improves the numerical accuracy of the results. In fact, a similar trick has been already known among some of the researchers (see the open-source code of [16, 4] for example). Specifically, we reindex the points, if necessary, to ensure that
    $m_{13}\leq m_{12}\leq m_{23}.$ (36)
    Note that this is different from [16, 4] where the indexing is based on the relative magnitudes of $s_{12}$, $s_{13}$ and $s_{23}$.
  3. 3.
    Finding $\mathbf{d}_{\mathbf{3}}$: Once $x$ and $y$ values are found, we can choose to use either (14), (15) or (16) to compute $d_{3}$. Theoretically, the choice of equation should not matter. In practice, however, we found that using (16) achieves higher numerical accuracy than using the other two.
Our method is summarized in Alg. 1.
Algorithm 1 Proposed P3P method
Input: 3D points, $\mathbf{X}_{i}$, and the corresponding unit bearing vectors, $\mathbf{m}_{i}$, $i=1,2,3$.
Output: Poses $\{\mathbf{R}_{j},\mathbf{t}_{j}\}_{j=1,\dots,N}$, where $N\leq 4.$
1:Compute $m_{12},m_{13},m_{23}$ using (4)–(6).
2:Reindex, if necessary, the points and bearing vectors so that (36) is satisfied.
3:Compute $s_{12},s_{13},s_{23}$ using (7)–(9).
4:Compute $c_{4},c_{3},c_{2},c_{1},c_{0}$ using (27)–(31).
5:if $\left|\frac{c_{3}}{c_{4}}\right|>10$ then
6:  Solve (26) using Ferrari-Lagrange method [21].
7:else
8:  Solve (26) using Classical Ferrari method [3].
9:end if
10:for every positive real root $x$ do
11:  Compute $y$ using (22)–(25).
12:  if $y>0$ then
13:   Compute $d_{3}$ using (16).
14:   Compute $d_{1}$ and $d_{2}$ using (13).
15:   Refine $d_{1},d_{2},d_{3}$ using Gauss-Newton method on (10)–(12).
16:   Compute $\mathbf{R}$ and $\mathbf{t}$ using (32)–(35).
17:  end if
18:end for
5 Results
In this section, we compare our method against several state-of-the-art solvers: namely, the ones by Wu et al. [23], Ding et al. [4] [^3] We test both the original version, Ding_old, published in November 2023, and the updated version, Ding_new, published in April 2024. For both versions, we use the ‘Adjoint’ method, as recommended by the authors., Persson and Nordberg [17], Nakano [16] [^4] The author of [16] published only the MATLAB code. To ensure a fair comparison, we use the C++ implementation provided by the authors of [4] at https://github.com/yaqding/P3P (as in [23])., Ke and Roumeliotis [11], and Kneip et al. [12]. All methods are implemented in C++ and run on a laptop CPU with i7-7700HQ 2.8GHz. Our evaluation strategy is aligned with the approaches used in [17, 4, 23]. First, the synthetic dataset is generated as follows:
  1. 1.
    Generate three random image points with coordinates $(u_{i},v_{i})$ such that $u_{i},v_{i}\sim\mathcal{U}(-1,1)$ for $i=1,2,3$.
  2. 2.
    Generate the unit bearing vectors, $\mathbf{m}_{i}$:
    $\mathbf{m}_{i}=\frac{[u_{i},v_{i},1]^{\top}}{\sqrt{u_{i}^{2}+v_{i}^{2}+1}}.$ (37)
  3. 3.
    Generate the corresponding 3D points $\mathbf{X}_{i}$ with a random depth $d_{i}\sim\mathcal{U}(0.1,10)$:
    $\mathbf{X}_{i}=d_{i}\mathbf{m}_{i}.$ (38)
  4. 4.
    Generate a random ground-truth rotation matrix $\mathbf{R}_{\text{gt}}$ and a translation vector of length 111, $\mathbf{t}_{\text{gt}}$.
  5. 5.
    Apply this random rigid body transformation to the points:
    $\mathbf{X}_{i}\leftarrow\mathbf{R}_{\text{gt}}^{\top}(\mathbf{X}_{i}-\mathbf{t}_{i}).$ (39)
Next, we evaluate each method on the same dataset and report the pose error used in [4, 23]:
$\xi=\lVert\mathbf{R}_{\text{gt}}-\mathbf{R}_{\text{est}}\rVert_{L1}+\lVert\mathbf{t}_{\text{gt}}-\mathbf{t}_{\text{est}}\rVert_{L1}.$ (40)
Based on this error metric, we determine how many solutions fall into each of the following categories (as in [4, 23]):
  1. 1.
    Valid solutions: All solutions $(\mathbf{R},\mathbf{t})$ returned by the solver.
  2. 2.
    Unique solutions: Solutions that have a valid rotation matrix [^5] We consider a rotation matrix to be valid if it meets the following conditions: (1) $|\det{(\mathbf{R})-1}|<10^{-6}$, (2) $\lVert\mathbf{R}^{\top}\mathbf{R}-\mathbf{I}\rVert_{L1}<10^{-6}$, and (3) $|1-\text{Quaternion norm of }\mathbf{R}|<10^{-5}$. and a reprojection error in the normalized image coordinates below $10^{-4}$, after removing the duplicates.
  3. 3.
    Duplicates: When two solutions $(\mathbf{R}_{1},\mathbf{t}_{1})$ and $(\mathbf{R}_{2},\mathbf{t}_{2})$ satisfy $\lVert\mathbf{R}_{1}-\mathbf{R}_{2}\rVert_{L1}+\lVert\mathbf{t}_{1}-\mathbf{t}_{2}\rVert_{L1}<10^{-5}$, we say that we have a duplicate.
  4. 4.
    Good solutions: The number of simulated problems where at least one unique solution is found.
  5. 5.
    No solution: The number of simulated problems where no unique solution is found.
  6. 6.
    Ground truth: The number of simulated problems where at least one solution satisfies $\xi<10^{-6}.$
  7. 7.
    Incorrect: The number of valid solutions (including the duplicates) that do not meet the criteria for unique solutions.
The main results are presented in Tab. 1. It can be seen that Ding_new [4] is the most robust method. In terms of the number of ground-truth solutions obtained, our method is the second best: The difference between ours and Ding_new [4] is only 888 out of $10^{8}$ problems, amounting to a statistically negligible difference of 0.000008% of the total problems. In terms of the number of good solutions, the difference is only 4 out of $10^{8}$, which amounts to 0.000004%. We also found that the method by Wu et al. [23] performs slightly worse than the authors reported: In [23], it achieves the ground truth in 99.99997% of cases, while in our evaluation, it does so in 99.99989% of cases. We speculate that two potential causes may account for this discrepancy. First, we use a sample size that is 10 times larger than that used in [23], so there is a lower risk of sampling bias. Second, the authors of [23] might have mistakenly simulated the depths $d_{i}$ in (38) using a uniform distribution within the interval [0.1,100][0.1,100][ 0.1 , 100 ], instead of [0.1,10][0.1,10][ 0.1 , 10 ] as stated in their paper [^6] We suspect this for two main reasons: First, the authors’ public source code uses the distribution $\mathcal{U}(0.1,100)$ for the depths, and the file containing this line is identical to one uploaded by the authors of [4] where the same distribution is used. Second, our experiment confirmed that using the distribution $\mathcal{U}(0.1,100)$ for the depths improves the results of their method, producing numbers very close to those reported in [23]..
Table 2 compares the mean, median and maximum errors of the ground-truth solutions produced by each method. Looking at the mean errors, Ding_new [4] once again demonstrates the highest numerical stability, with our method ranking a close second. With an average numerical difference of just $1.8\times 10^{-13}$, the two methods exhibit a comparable level of numerical stability.
Table 3 compares the running times of each method. Overall, the two fastest methods are the one by Wu et al. [23] and ours. Looking at the mean times, the method by Wu et al. [23] is faster than ours, but only by 6.6 nanoseconds on average, a difference that is practically negligible.
Finally, in Tab. 4, we conduct an ablation study to evaluate the impact of the heuristics described in Section 4.2. The result illustrates that even seemingly minor details can have a non-negligible impact in practice.
Method Ours Wu [23] Ding_new [4] Ding_old [4] Persson [17] Ke [11] Kneip [12] Nakano [16]
Valid 168853631 168853618 168853640 168853633 168853625 174550801 242373408 196891118
Unique 168853586 168853563 168853619 168853604 168853528 168866306 168876463 168844644
Duplicates ($\downarrow$) 22 18 17 16 20 1893830 29666 108
Good ($\uparrow$) 99999995 99999984 $99$ 99999996 99999969 99997580 99997085 99994088
No solution ($\downarrow$) 5 16 1 4 31 2420 2915 5912
Ground truth ($\uparrow$) 99999979 99999890 99999987 99999933 99999777 99989030 99957466 99993317
Incorrect ($\downarrow$) 23 37 4 11 77 3790665 73472739 28046366
Table 1: Accuracy comparison of the eight solvers on $10^{8}$ simulated problems. The symbol $\uparrow$ indicates that the higher the better. The symbol $\downarrow$ indicates that the lower the better. The best results are highlighted in bold.
Method Mean Median Maximum
Ours 1.514e-12 3.558e-14 9.662e-7
Wu [23] 6.918e-12 3.266e-14 9.666e-7
Ding_new [4] 1.335e-12 2.266e-14 8.306e-7
Ding_old [4] 4.647e-12 3.010e-14 9.925e-7
Persson [17] 1.082e-11 4.216e-14 9.960e-7
Ke [11] 9.195e-11 1.835e-14 1.000e-6
Kneip [12] 2.756e-10 4.317e-14 1.000e-6
Nakano [16] 1.694e-12 2.526e-14 9.589e-7
Table 2: Comparison of the mean, median and maximum errors of the ground-truth solutions returned by each method (given in the second row from the bottom of Tab. 1).
Time (ns)
Ours
Wu [23]
Ding_new [4]
Ding_old [4]
Persson [17]
Ke [11]
Kneip [12]
Nakano [16]
Mean 254.7 248.1 257.2 267.8 305.9 515.1 826.5 1071.7
Median 256.1 248.0 256.6 270.5 304.3 514.8 831.9 1075.7
Minimum 237.2 228.5 236.5 247.7 286.6 467.1 747.2 954.9
Maximum 280.2 290.9 292.4 330.1 345.1 608.7 937.8 1434.8
Table 3: Time comparison of the eight solvers on $10^{8}$ simulated problems: For each individual problem, we repeat the test 10 times and average the measurements. Then, we report the mean, median, minimum and maximum of these $10^{8}$ average results.
Method Ours (Baseline) Using only Ferrari- Lagrange method [21] Using only Classical Ferrari method [3] Without reindexing Using (14) instead of (16) Using (15) instead of (16)
Valid 168853631 168853016 168853148 168853457 168853631 168853631
Unique 168853586 168852752 168852972 168853340 168853574 168853574
Duplicates ($\downarrow$) 22 84 50 41 21 21
Good ($\uparrow$) 99999995 99999700 99999867 99999951 99999992 99999992
No solution ($\downarrow$) 5 300 133 49 8 8
Ground truth ($\uparrow$) 99999979 99999600 99999645 99999836 99999948 99999946
Incorrect ($\downarrow$) 23 180 126 76 36 36
Table 4: Ablation study on $10^{8}$ simulated problems: We evaluate the impact of the heuristics described in Section 4.2.
6 Discussion
A key strength of our method lies in its simplicity. By reducing the P3P problem to solving a single quartic polynomial with straightforward coefficients, we avoid the complexity often seen in other approaches. Despite its simplicity, our method achieves accuracy and runtime on par with state-of-the-art solvers. In many cases, it even outperforms more elaborate methods in terms of stability and robustness. This demonstrates that a well-designed minimal formulation can match or even outperform more sophisticated alternatives.
However, the quartic root-finding step remains the most numerically sensitive component of the pipeline. Our adaptive choice between two Ferrari-based solvers helps mitigate this, but certain failure cases remain where existing methods like Ding_new [4] offer marginally better numerical stability. Exploring more robust root-finding strategies or methods that exploit the structure of our specific quartic could further enhance reliability and efficiency. We leave this for future work.
7 Conclusion
In this work, we revisited the P3P problem and showed that it can be solved by finding the roots of a quartic polynomial with compact, analytically simple coefficients. We also introduced several implementation strategies that significantly improve numerical stability in practice. Extensive evaluation demonstrated that our method delivers accuracy and runtime performance on par with state-of-the-art approaches. Beyond its empirical performance, the algebraic simplicity of our method makes it a valuable tool for teaching geometric vision concepts, developing lightweight systems for edge computing, and providing a clear baseline for future theoretical work on minimal solvers. Our code is publicly available at https://github.com/sunghoon031/P3P.